OpenStudy (anonymous):

Hi, can someone please help me with this problem? 4 8/9 + 6 1/5 + 2 4/5 =

OpenStudy (amigatour):

That's adding fractions. First step, find their common denominators then add.

OpenStudy (anonymous):

That would be 40. Then what is the next step?

OpenStudy (anonymous):

Is it then do you make a improper fraction?

OpenStudy (amigatour):

Yes

OpenStudy (amigatour):

Ok. I;m having problem with my screen. It's hanging up.. I'm going to do it on paper and post here.

OpenStudy (anonymous):

So I have 168/40 + 25/40+ 84/40=

OpenStudy (anonymous):

277/40?

OpenStudy (anonymous):

How did you get your equation? Thank you for your help!

OpenStudy (ranga):

how is 40 the common denominator?

OpenStudy (ranga):

4 8/9 + 6 1/5 + 2 4/5 you can add the whole numbers separately and the fractions separately: 4 + 6 + 2 + 8/9 + 1/5 + 4/5 12 + 8/9 + 1/5 + 4/5 add the fractions using common denominator.

OpenStudy (ranga):

4/5 + 1/5 has 5 as the common denominator and they can be added up: 4/5 + 1/5 = (4 + 1)/5 = 5/5 = 1

OpenStudy (anonymous):

I am just learning how to do this in school and don't understand.

OpenStudy (ranga):

which part you are having trouble with?

OpenStudy (amigatour):

It's mixed fractions, right?

OpenStudy (anonymous):

How to get the common denominator?

OpenStudy (anonymous):

Yes mixed fractions.

OpenStudy (amigatour):

With mixed fractions, you have to make them improper fractions-if their denominator is different otherwise... Check the photo I have posted.

OpenStudy (anonymous):

Ok.

OpenStudy (anonymous):

Thank you, showing it step by step really helped! Thank you for your time!

OpenStudy (amigatour):

Yw

OpenStudy (ranga):

Here is another way to do the same problem: 4 8/9 + 6 1/5 + 2 4/5 you can add the whole numbers separately and the fractions separately: 4 + 6 + 2 + 8/9 + 1/5 + 4/5 12 + 8/9 + 1/5 + 4/5 12 + 8/9 + (1 + 4)/5 12 + 8/9 + 1 13 + 8/9 13 8/9

OpenStudy (anonymous):

Thank you Ranga!

OpenStudy (ranga):

you are welcome.